Wildhorse Company manufactures one product. On December 31, 2019, Wildhorse adopted the dollar-value LIFO inventory method. TI don't think the first answer is correct, please type the other three thank u
Answer #1

Calculate ending inventory

Ending inventory
2020 1272600/1.05 = 1212000 1050000+162000*1.05 = 1220100
2021 1784800/1.15 = 1552000 1220100+(340000*1.15) = 1611100
2022 1915000/1.25 = 1532000 1220100+(320000*1.15) = 1588100